Parker Plastic, Inc., manufactures plastic mats to use with rolling office chairs. Its standard cost information for last year follows:

Standard Quantity Standard Price (Rate) Standard Unit Cost
Direct materials (plastic) 10 sq ft. $ 0.82 per sq. ft. $ 8.20
Direct labor 0.25 hr. $ 10.20 per hr. 2.55
Variable manufacturing overhead (based on direct labor hours) 0.25 hr. $ 1.40 per hr. 0.35
Fixed manufacturing overhead $270,600 ÷ 902,000 units) 0.30


Parker Plastic had the following actual results for the past year:

Number of units produced and sold 1,020,000
Number of square feet of plastic used 11,000,000
Cost of plastic purchased and used $ 8,800,000
Number of labor hours worked 300,000
Direct labor cost $ 3,000,000
Variable overhead cost $ 400,000
Fixed overhead cost $ 357,000


Required:
Calculate Parker Plastic’s variable overhead rate and efficiency variances and its over- or underapplied variable overhead. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for Favorable/Overapplied and "U" for Unfavorable/Underapplied.)

Variable overhead rate variance

Variable overhead efficiency variance

Variable overhead spending variance
Answer #1

Voh rate variance = (SR-AR) *AH

= (1.40-400,000/300,000)*300,000

= 20,000 F

Voh efficiency variance = (SH-AH) *SR

= (1,020,000*0.25 - 300,000)*1.40

= 63,000 U

Voh spending variance = 20,000 F + 63,000 U

= 43,000 U